OpenStudy (anonymous):

A printer printed 68 letters in 4 seconds. If it prints letters at a constant speed, it can print ______ letters in 30 seconds.

OpenStudy (anonymous):

First you do, 68/4 to give you the amount of letters it can print in 1 second. This is 17. Then you do 17*30= 510. I think this is it...

OpenStudy (texaschic101):

proportion time... 68/4 = x/30 --- 68 letters in 4 sec = x letters in 30 sec cross multiply 4x = (68)(30) 4x = 2040 -- divide both sides by 4 x = 510 510 letters in 30 sec

OpenStudy (anonymous):

Oh yes I was righttttttttt ^-^ Hehehe